This is a problem I missed on a quiz. Please, can anyone help me work this problem? Here it is: A small area in front of a building is triangular in shape. The perimeter is 49 meters. The second side is one-fourth of the first side in length. The third side is 4 meters more than the first side. FInd the length in meters of each side of the triangular region.
Answer by nerdybill(7384)

.
Let x = length of first side
then from "The second side is one-fourth of the first side"
x/4 = length of second side
and from "third side is 4 meters more than the first side"
x+4 = length of third side
.
x + x/4 + (x+4) = 49
Multiplying both sides by 4:
4x + x + 4x + 16 = 196
9x + 16 = 196
9x = 180
x = 20 meters (first side)
.
second side:
x/4 = 20/4 = 5 meters (second side)
.
third side:
x+4 = 20+4 = 24 meters (third side)